OpenStudy (anonymous):

Simplify √-121 √ = radical symbol

hartnn (hartnn):

as told earlier : \(\sqrt{-1}=i\) so, \(\sqrt{-121}=\sqrt{-1}\sqrt{121}=..?\)

OpenStudy (anonymous):

Where did the 'i' come in at ?!?!

hartnn (hartnn):

i is actually the square root of negative 1, so i slpit square root of negative 1 in to 2 parts, one having 'i'.

OpenStudy (anonymous):

My choices are a. -11i b.11i c.-11 d. 11

hartnn (hartnn):

*split square root of -121...

OpenStudy (anonymous):

Is th answer 11

hartnn (hartnn):

ok, out of the 2 parts here \(\sqrt{-121}=\sqrt{-1}\sqrt{121}=..?\), 1st part equals 'i' what does 2nd part square root of 121 equal ?? and no.

OpenStudy (anonymous):

I don't know how to do square root.... I don't have a calc and I just want the answer :)

hartnn (hartnn):

but you have google :) type in square root of 121.....

OpenStudy (anonymous):

It told me 11..

hartnn (hartnn):

thats correct. so \(\sqrt{-121}= i \times 11\)

OpenStudy (anonymous):

soooo 11i ?

hartnn (hartnn):

yup.

OpenStudy (anonymous):

What aboutttt \[\sqrt{-72}\]

hartnn (hartnn):

same way, \(\sqrt{-72}=\sqrt{-1}\sqrt{72}=i \times \sqrt{72}\) you just need to simplify, square root of 72

OpenStudy (anonymous):

8.48528137424....

hartnn (hartnn):

but i think your choices will have radicals in them.

OpenStudy (anonymous):

My choices are a. \[-6\sqrt{2}\] b. \[6\sqrt{-2}\] c.\[6\sqrt{2i}\] d.\[6i \sqrt{2}\]

hartnn (hartnn):

can you factor 72 ?

OpenStudy (anonymous):

I don't know how to factor.. Isn't that where you find 2 of the GCF ?

hartnn (hartnn):

72 = 2*36 and whats the square root of 36 ?